Medal Corporation uses direct labor hours to allocate factory overhead. During the period, the company produced 5,100 units and incurred variable factory overhead costs of $21,000, Use the information below to determine the following (round dollars to the nearest cent): Medal Corporation Factory Overhead Cost Budget For the three months ending March 31, 2015 100% I 10% Percent of normal capacity Units Direct labor hours (0.75 hr. per unit Budgeted factory overhead: 90% ced 5,000 3,750 5,500 4,125 4 Variable cost: Indirect factory labor Indirect materials Utilities 8,100 9,000 3,600 18,450 12,000 0,000 4,000 20,500 12,000 0,000 Total variable cost 22,5 Fixed costs Su sor salaries 12,000 Total fixed costs Total fac overhead cost 38.45 a. Variable factory overhead rate b. Variable factory overhead controllable variance c. Is the variable factory overhead controllable variance favorable or unfavorable? d. Fixed factory overhead rate e. Fixed factory overhead volume variance f. Is the fixed factory overhead volume variance favorable or unfavorable? g. Determine the total factory overhead cost variance by comparing actual and applied factory overhead. Is the variance favorable or unfavorable?
Answer #1

Answer to Question a:

x1=Variable Cost at 1` 00% Capacity =$20500 x2=Variable Cost at given Capacity= $21000

y1= No of units at 100% Capacity= 5000units y2= No of units at given Capacity= 5100 units

Variable Factory Overhead Rate=

Change in Variable Cost/ Change in Units= x2-x1/y2-y1 = $21000- $20500/ (5100-5000)

=$5 per unit

Answer to Question b:

Budgeted OH Rate = (40500- 38450)/(5000-4500)

= $4.1

Variable Factory Overhead Controllable Variance=

Actual Variable Factory Overhead- Budgeted Variable Factory Overhead=

= $21000- ( 5100 units * $4.1)

= $21000- $20910

=$90(A)

Answer to Question c:

Variable Factory Overhead Controllable Variance is an adverse variance as actual Variable Factory Overhead is more than budgeted variable Factory Overhead.

Answer to Question d:

Fixed Factory Overhead Rate =

= (40500- 38450)/(5000-4500)

= $4.1

Answer to Question e:

Budgeted Fixed Overhead=$20000

Total Applied Fixed Overhead at 5100 units = 5100 units * $4.1 = $20910

Fixed Factory Overhead Volume Variance= Applied Fixed Overhead- Budgeted Fixed Overhead

= $20910-$20000

=$910(F)

  

Answer to Question f:

Fixed Factory Overhead Volume Variance is Favourable as OH absorbed is more than budgeted.

Answer to Question g

Actual Total Factory Overhead Cost at 5100 units = Variable Factory Overhead Cost+Total Fixed Costs

=$21000+ $20000 = $41000

Applied Total Factory Overhead Cost at 5100 units = ($5100*4.1+$20000)

= $40910

Total Factory Overhead Cost Variance= Actual Cost - Applied Cost

= $41000- $40910

=$90(A)

As actual  Factory Overhead Cost is more than recovered it is an adverse Variance.